# Large Language Models are In-Context Molecule Learners

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional performance in biochemical tasks, especially the molecule caption translation task, which aims to bridge the gap between molecules and natural language texts. However, previous methods in adapting LLMs to the molecule-caption translation task required extra domain-specific pre-training stages, suffered weak alignment between molecular and textual spaces, or imposed stringent demands on the scale of LLMs. To resolve the challenges, we propose In-Context Molecule Adaptation (ICMA), as a new paradigm allowing LLMs to learn the molecule-text alignment from context examples via In-Context Molecule Tuning. Specifically, ICMA incorporates the following three stages: Hybrid Context Retrieval, Post-retrieval Re-ranking, and In-context Molecule Tuning. Initially, Hybrid Context Retrieval utilizes BM25 Caption Retrieval and Molecule Graph Retrieval to retrieve informative context examples. Additionally, we also propose Post-retrieval Re-ranking with Sequence Reversal and Random Walk to further improve the quality of retrieval results. Finally, In-Context Molecule Tuning unlocks the in-context molecule learning capability of LLMs with retrieved examples and adapts the parameters of LLMs for the molecule-caption translation task. Experimental results demonstrate that ICMA can empower LLMs to achieve state-of-the-art or comparable performance without extra training corpora and intricate structures, showing that LLMs are inherently in-context molecule learners.
1 Introduction
Molecules play a crucial role across various fields, such as medicine  Ding et al. (2019), agriculture  Twyman et al. (2003), and material science  Higuchi et al. (2023), as they are widely used in the development of drugs, fertilizers, and advanced materials. Recently, LLMs have demonstrated remarkable success in the molecular domain, as molecules can be represented as Simplified Molecular-Input Line-Entry System (SMILES) strings  Weininger (1988), which can be comprehended and generated by LLMs in a similar manner to natural languages. To further bridge the gap between the molecules and natural languages,  Edwards et al. (2022) proposes the molecule-caption translation task, which comprises two sub-tasks: molecule captioning (Mol2Cap) and text-based de novo molecule generation (Cap2Mol). Specifically, Mol2Cap involves generating a textual description that elucidates the features of the given molecule, while Cap2Mol focuses on predicting the exact molecule based on the textual caption. The study of the molecule-caption translation task offers an accessible and chemist-friendly venue for molecule discovery, which has raised wide research focus.
@@ 1 Introduction
Generally, there are two main paradigms for adapting LLMs to the molecule-caption retrieval task. The first paradigm is the domain-specific pre-training & fine-tuning. For instance, MolT5  Edwards et al. (2022) first proposes and handles the molecule-caption translation task as the language translation task, pre-training the MolT5 model with chemical corpora like PubChem  Kim et al. (2019) and then fine-tuning the model on the ChEBI-20 dataset  Edwards et al. (2021). Additionally, MoMu  Su et al. (2022) and MolCA  Liu et al. (2023b) introduce an extra modality alignment stage before fine-tuning on downstream tasks, which aligns the output of 2D molecule graph encoder with the input space of LLMs. In contrast, the other paradigm involves prompting and utilizing the in-context learning capability of LLMs. For example, MolReGPT  Li et al. (2023) introduces In-Context Few-Shot Molecule Learning, prompting general LLMs like GPT-3.5 and GPT-4 to achieve competitive performance without extra parameter adaptations.
However, the current paradigms still face challenges. On one hand, the domain-specific pre-training & fine-tuning paradigm requires extra pre-training stages (i.e., domain-specific pre-training and modality alignment), which is challenging due to the scarcity of high-quality chemical datasets, especially molecule-caption pairs, making this paradigm difficult to scale up to the most advanced LLMs with billion parameters. Besides, the domain-specific pre-training & fine-tuning paradigm also suffers from weak alignment between molecules and texts, as phrases in molecule captions often indicate specific sub-structures of molecules rather than the entire molecule. Despite attempts to introduce extra modalities for better alignment Su et al. (2022); Liu et al. (2023b), the integration of the additional modalities (e.g., 2D molecule graph) is still focused on the entire graph level and can only be applied to the molecule captioning task, while ignoring the text-based generation of molecules, which is much more valuable for drug discovery. On the other hand, the in-context learning & prompting paradigm puts a harsh requirement on LLMs’ emergent capabilities, such as reasoning and in-context learning abilities. However, LLMs with these emergent capabilities usually have billions of parameters, making them computationally expensive. Consequently, there is a demand for a unified and efficient approach that effectively enhances the performance of the most advanced LLMs in both two sub-tasks of molecule-caption translation.
@@
Specifically, ICMA incorporates three stages: Hybrid Context Retrieval, Post-retrieval Re-ranking, and In-context Molecule Tuning. In the initial stage, Hybrid Context Retrieval, we employ Caption Retrieval and Molecule Graph Retrieval to fetch similar captions and molecules, respectively. Subsequently, we introduce the Post-retrieval Re-ranking stage to enhance the quality of the retrieval algorithms. This stage incorporates two innovative strategies: Sequence Reversal and Random Walk, which aim to refine and reprioritize the retrieved examples. Finally, we apply In-context Molecule Tuning to adapt the parameters of LLMs, enabling them to learn from the contextual mappings and effectively utilize the learned knowledge for the current generation task. Experiments are conducted across two real-world molecule-caption translation datasets, ChEBI-20 and PubChem324k. Results show that ICMA could enable LLMs to achieve state-of-the-art or comparable performance in both the two sub-tasks (i.e., Mol2Cap and Cap2Mol). Meanwhile, we also study the factors related to the model performance, including retrieval algorithms, context settings (i.e., context example number and maximum input length), and model scales. Lastly, ablation studies are conducted to justify the effectiveness of Post-retrieval Re-ranking components.
@@
  • We conduct synthetic experiments, and the results show that our method enables LLMs to outperform previous paradigms, enabling better alignment between molecules and texts. Notably, our approach elevates the Mistral-7b model to establish state-of-the-art or comparable performance across both the two sub-tasks of molecule-caption translation, achieving 0.581 BLEU-4 score in Mol2Cap and 0.460 exact-matched score in Cap2Mol. Additionally, we comprehensively study the mechanism and influential factors of ICMA, showing that LLMs are inherently in-context molecule learners.
2 Related Work
@@
2.1 Molecule-Caption Translation
@@
Inspired by the image captioning task, Edwards et al. (2021) introduces a new dataset, ChEBI-20, with pairs of molecules and manually labeled captions that describe the molecular properties. The molecule-caption translation task was initially proposed in MolT5 Edwards et al. (2022). Meanwhile, MolT5 proposes a T5 model that is jointly pre-trained on molecule SMILES and general text corpus. MolXPT  Liu et al. (2023a) pre-trains a GPT model by introducing extra-wrapped texts as the pre-training corpus, demonstrating better molecule-text alignment. However, the generation of SMILES strings suffers from the problem of invalid SMILES due to the mismatches of brackets. To overcome the generation issue of SMILES strings, BioT5  Pei et al. (2023) introduces Self-referencing Embedded Strings (SELFIES)  Krenn et al. (2020) instead of SMILES strings to represent molecules in LLMs and proposes the BioT5 model that is jointly trained on single-modal data, wrapped text data, and molecule/protein-description pairs. Meanwhile, as molecules can also be represented as graphs, some methods focus on molecule understanding by introducing molecule graph information. For example, MoMu  Su et al. (2022) first proposes a graph encoder to encode molecule graph information and utilizes contrastive learning to bridge the semantic gap between the graph encoder and the LLM. To better fuse the molecule graph information, MolCA  Liu et al. (2023b) follows the BLIP-2 and utilizes a Q-Former to project the output of the graph encoder into the LLMs, showing better molecule understanding performance. However, most of these methods still adhere to the pre-training & fine-tuning paradigm, which necessitates the joint pre-training of LLMs on both general text and extra chemical domain corpora. As the size of the training corpora and model weights of LLMs continue to increase, this approach has become extremely inefficient. To address this issue, MolReGPT  Li et al. (2023) proposes the In-Context Few-Shot Molecule Learning to enable LLMs to learn the molecule-caption translation task from the context examples without modifying the model weights while still achieving comparable performance to these fine-tuned methods.
2.2 In-Context Learning
@@
With the scaling of model size and corpus size  Brown et al. (2020); Chowdhery et al. (2023), LLMs emerge the in-context learning capability  Wang et al. (2022), which enables LLMs to learn from contexts augmented with several examples  Dong et al. (2022). By utilizing the capability of ICL, LLMs can solve complex tasks without the necessity of being fine-tuned. For instance, with a few examples, GPT-3 could demonstrate similar performance to fine-tuned models in unseen tasks  Brown et al. (2020). What’s more, based on context examples, LLMs could achieve better mathematical reasoning ability with the assistance of chain-of-thought (CoT)  Wei et al. (2022). With the powerful in-context learning capabilities,  Edwards et al. (2023) proposes in-context drug synergy learning to apply LLMs for personalized drug synergy prediction and drug design, while  Jablonka et al. (2024) fine-tunes LLMs such as GPT-3 with chemical questions and answers to solve predictive tasks In the scenario of molecule-caption translation, MolReGPT  Li et al. (2023) proves the retrieval quality is closely related to the model performance, while it still imposes strict requirements on the model scale, as the in-context learning capability only becomes apparent when the model weights reach a certain size.
3 In-Context Molecule Adaptation
@@
In this section, we introduce In-Context Molecule Adaptation (ICMA) as a novel paradigm to adapt LLMs to molecule-caption translation. As shown in Figure 2, ICMA incorporates three stages, including Hybrid Context Retrieval, Post-retrieval Re-ranking, and In-context Molecule Tuning. Specifically, Hybrid Context Retrieval first retrieves $N$ rough examples from the training set $\mathcal{D}$ by calculating the similarity between the current query and molecule-caption pairs. After that, Post-retrieval Re-ranking with Random Walk and Sequence Reversal is adopted to obtain $n$ refined examples from the $N$ rough examples. Finally, the In-context Molecule Tuning can be performed to update the parameters of LLMs to learn the molecule-text alignment from refined context examples.
3.1 Hybrid Context Retrieval
@@
For the Mol2Cap task, ICMA adopts Molecule Graph Retrieval to better refine the retrieval quality. Previously, MolReGPT  Li et al. (2023) utilizes the Morgan Fingerprints (Morgan FTS) and Dice similarity to evaluate the similarity between molecules, which encodes the pre-defined handcraft structures as the embedding of the molecule. However, Morgan FTS are typically based on handcrafted chemical feature extraction, which may have limited capability to capture the comprehensive information of complex molecule structures and properties. On the other side, Graph Neural Networks (GNNs), pre-trained on millions of molecules, could effectively capture topological structures of molecule graphs, providing complete chemical semantics Xia et al. (2023). This makes GNNs a better option for molecule similarity calculation. In ICMA, we adopt a pre-trained GNN encoder to obtain the molecule graph embeddings:
@@
where $\textbf{e}_{m}$ denotes the embedding of the given 2D graph $g_{m}$ of the molecule $m$. Specifically, we adopt Mole-BERT  Xia et al. (2022) as the GNN encoder.
@@
In the Cap2Mol task, we inherit the BM25 Caption Retrieval  Robertson et al. (2009) from MolReGPT  Li et al. (2023) as it focuses on the detail matching of molecule captions, showing competitive performance and is much faster than LLM-based methods like Sentencebert  Reimers and Gurevych (2019). Specifically, given captions in the test set as query captions $Q_{c}$ and the training set of captions $\mathcal{D}_{c}$, the caption similarity ranking function $\mathcal{R}^{c}$ can be denoted as:
@@
3.2 Post-retrieval Re-ranking
Although refined retrieval algorithms could bring better retrieval quality, there are still some problems considering the arrangement of context examples. Thus, we propose Post-retrieval Re-ranking with Random Walk and Sequence Reversal to re-rank the priorities and positions of context examples, thus enhancing In-Context Molecule Tuning quality.
3.2.1 Random Walk
The molecules ranked top by retrieval algorithms can sometimes share too many overlaps, impairing the informativeness of context examples. In this case, for the context diversity and generalization performance of ICMA, it is necessary to give these less similar examples a chance to be visited. Inspired by the random walk mechanism in graph theory, we propose random walk as a post-retrieval method to select examples from the top-$N$ retrieved results so that examples with lower rank still have a chance to be selected, which provides more useful information and complements the context diversity. Mathematically, we adopt a dynamic chance for examples with different ranks. Specifically, for the $j$-th example in the $N$ rough results, where $1\leq j\leq$N, the possibility of skipping is represented as $(9*\frac{N-j}{N-1})\%$, which reduces as the rank moves down to ensure that at least one example is selected for the context. If the $j$-th example is skipped, then it will perform a Random Walk in the remaining examples until an example is selected. Practically, if $N$ is too large (i.e., $N\gg n$), the retrieval quality will be impaired, while if $N$ is too small (i.e., $N\approx n$), it might hurt the diversity. Therefore, considering $n$ is not quite large and usually less than 5 due to the context length limitation, $N$ is set to 10 in this work.
3.2.2 Sequence Reversal
Due to the training strategy and the inherent characteristic of natural languages, LLMs have difficulty capturing long-range dependencies or relationships between words that are far apart in the input text, namely the distance dependency. The positions of examples in the context might influence the generation results due to the distance dependency of LLMs  Frermann et al. (2023). In this case, it is of significance to put the most informative example exactly near the current input query.

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional performance in biochemical tasks, especially the molecule caption translation task, which aims to bridge the gap between molecules and natural language texts. However, previous methods in adapting LLMs to the molecule-caption translation task required extra domain-specific pre-training stages, suffered weak alignment between molecular and textual spaces, or imposed stringent demands on the scale of LLMs. To resolve the challenges, we propose In-Context Molecule Adaptation (ICMA), as a new paradigm allowing LLMs to learn the molecule-text alignment from context examples via In-Context Molecule Tuning. Specifically, ICMA incorporates the following three stages: Hybrid Context Retrieval, Post-retrieval Re-ranking, and In-context Molecule Tuning. Initially, Hybrid Context Retrieval utilizes BM25 Caption Retrieval and Molecule Graph Retrieval to retrieve similar informative context examples. Additionally, Post-retrieval Re-ranking is composed of Sequence Reversal and Random Walk selection to further improve the quality of retrieval results. Finally, In-Context Molecule Tuning unlocks the in-context learning and reasoning capability of LLMs with the retrieved examples and adapts the parameters of LLMs for better alignment between molecules and texts. Experimental results demonstrate that ICMA can empower LLMs to achieve state-of-the-art or comparable performance without extra training corpora and intricate structures, showing that LLMs are inherently in-context molecule learners.
Index Terms:
Drug Discovery, Large Language Models (LLMs), In-context Tuning, Retrieval Augmented Generation.
I Introduction
Molecules play a crucial role across various fields, such as medicine [1], agriculture [2], and material science [3], as they are widely used in the development of drugs, fertilizers, and advanced materials. Recently, LLMs have demonstrated remarkable success in the molecular domain, as molecules can be represented as Simplified Molecular-Input Line-Entry System (SMILES) strings [4], which can be comprehended and generated by LLMs in a similar manner to natural languages. To further bridge the gap between the molecules and natural languages, MolT5 [5] proposes the molecule-caption translation task, which comprises two sub-tasks: molecule captioning (Mol2Cap) and text-based de novo molecule generation (Cap2Mol). Specifically, Mol2Cap involves generating a textual description that elucidates the features of the given molecule, while Cap2Mol focuses on predicting the exact molecule based on the textual caption. The study of the molecule-caption translation task offers an accessible and chemist-friendly venue for molecule discovery, which has raised wide research focus.
@@ I Introduction
Generally, there are two main paradigms for adapting LLMs to the molecule-caption retrieval task. The first paradigm is the domain-specific pre-training & fine-tuning. For instance, MolT5 [5] first proposes and handles the molecule-caption translation task as the language translation task, pre-training the MolT5 model with chemical corpora like PubChem [6] and then fine-tuning the model on the ChEBI-20 dataset [7]. Additionally, MoMu [8] and MolCA [9] introduce an extra modality alignment stage before fine-tuning on downstream tasks, which aligns the output of 2D molecule graph encoder with the input space of LLMs. In contrast, the other paradigm involves prompting and utilizing the in-context learning capability of LLMs. For example, MolReGPT [10] introduces In-Context Few-Shot Molecule Learning, prompting general LLMs like GPT-3.5 and GPT-4 to achieve competitive performance without extra parameter adaptations.
However, the current paradigms face critical challenges. On one hand, the domain-specific pre-training & fine-tuning paradigm requires extra pre-training stages (i.e., domain-specific pre-training and modality alignment), which is challenging due to the scarcity of high-quality chemical datasets, especially molecule-caption pairs, making this paradigm infeasible to scale up to the most advanced LLMs with billion parameters. Besides, the domain-specific pre-training & fine-tuning paradigm also suffers from weak alignment between molecules and texts, as phrases in molecule captions often indicate specific sub-structures of molecules rather than the entire molecule. Despite attempts to introduce extra modalities for better alignment [8, 9], the integration of the additional modalities (e.g., 2D molecule graph) is still focused on the entire graph level and can only be applied to the Mol2Cap task, while ignoring the text-based generation of molecules, which is much more valuable for drug discovery. On the other hand, the in-context learning & prompting paradigm puts a harsh requirement on LLMs’ emergent capabilities, such as reasoning and in-context learning abilities. However, LLMs with these emergent capabilities usually have billions of parameters, making them computationally expensive. Consequently, there is a demand for a unified and efficient approach that effectively enhances the performance of the most advanced LLMs in both two sub-tasks of molecule-caption translation.
@@
Specifically, ICMA incorporates three stages: Hybrid Context Retrieval, Post-retrieval Re-ranking, and In-context Molecule Tuning. In the initial stage, Hybrid Context Retrieval, we employ Caption Retrieval and Molecule Graph Retrieval to fetch similar captions and molecules, respectively. Subsequently, we introduce the Post-retrieval Re-ranking stage to enhance the quality of the retrieval algorithms. This stage incorporates two innovative strategies: Sequence Reversal and Random Walk, which aim to refine and reprioritize the retrieved examples. Finally, we apply In-context Molecule Tuning to adapt the parameters of LLMs, enabling them to learn from the contextual mappings and effectively ground the current generation task. Experiments are conducted across two real-world molecule-caption translation datasets, ChEBI-20 and PubChem324k. Results show that ICMA could enable LLMs to achieve state-of-the-art or comparable performance in both the two sub-tasks (i.e., Mol2Cap and Cap2Mol). Meanwhile, we also study the factors related to the model performance, including retrieval algorithms, context settings (i.e., context example number and maximum input length), model scales, and backbone LLMs. Lastly, the ablation study and detailed case study are conducted to justify the effectiveness of Post-retrieval Re-ranking components.
@@
  • We conduct synthetic experiments, and the results show that our method enables LLMs to outperform previous paradigms, enabling better alignment between molecules and texts. Notably, our approach elevates general LLM like Mistral-7B to establish superior performance across both the two sub-tasks of molecule-caption translation, achieving 0.581 BLEU-4 score in Mol2Cap and 0.460 exact-matched score in Cap2Mol. Additionally, we comprehensively study the mechanism and influential factors of ICMA, showing that LLMs are inherently in-context molecule learners.
II Related Work
@@
II-A Molecule-Caption Translation
@@
Inspired by the image captioning task, Edwards et al. [7] introduce a new dataset, ChEBI-20, with pairs of molecules and manually labeled captions that describe the molecular properties. The molecule-caption translation task was initially proposed in MolT5 [5]. Meanwhile, MolT5 proposes a T5 model that is jointly pre-trained on molecule SMILES and general text corpus. MolXPT [11] pre-trains a GPT model by introducing extra-wrapped texts as the pre-training corpus, demonstrating better molecule-text alignment. However, the generation of SMILES strings suffers from the problem of invalid SMILES due to the mismatches of brackets. To overcome the generation issue of SMILES strings, BioT5 [12] introduces Self-referencing Embedded Strings (SELFIES) [13] instead of SMILES strings to represent molecules in LLMs and proposes the BioT5 model that is jointly trained on single-modal data, wrapped text data, and molecule/protein-description pairs. Meanwhile, as molecules can also be represented as graphs, some methods focus on molecule understanding by introducing molecule graph information. For example, MoMu [8] first proposes a graph encoder to encode molecule graph information and utilizes contrastive learning to bridge the semantic gap between the graph encoder and the LLM. To better fuse the molecule graph information, MolCA [9] follows the BLIP-2 [14] and utilizes a Q-Former to project the output of the graph encoder into the LLMs, showing better molecule understanding performance. However, most of these methods still adhere to the pre-training & fine-tuning paradigm, which necessitates the joint pre-training of LLMs on both general text and extra chemical domain corpora. As the size of the training corpora and model weights of LLMs continue to increase, this approach has become extremely inefficient. To address this issue, MolReGPT [10] proposes the In-Context Few-Shot Molecule Learning to enable LLMs to learn the molecule-caption translation task from the context examples without modifying the model weights while still achieving comparable performance to these fine-tuned methods.
II-B In-Context Learning
@@
With the scaling of model size and corpus size [15, 16], LLMs emerge the in-context learning capability [17], which enables LLMs to learn from contexts augmented with several examples [18]. By utilizing the capability of ICL, LLMs can solve complex tasks without the necessity of being fine-tuned. For instance, with a few examples, GPT-3 could demonstrate similar performance to fine-tuned models in unseen tasks [15]. What’s more, based on context examples, LLMs could achieve better mathematical reasoning ability with the assistance of chain-of-thought (CoT) [19]. With the powerful in-context learning capabilities, Edwards et al. [20] propose in-context drug synergy learning to apply LLMs for personalized drug synergy prediction and drug design, while Jablonka et al. [21] fine-tunes LLMs such as GPT-3 with chemical questions and answers to solve predictive tasks In the scenario of molecule-caption translation, MolReGPT [10] proves the retrieval quality is closely related to the model performance, while it still imposes strict requirements on the model scale, as the in-context learning capability only becomes apparent when the model weights reach a certain size.
III In-Context Molecule Adaptation
@@
In this section, we introduce In-Context Molecule Adaptation (ICMA) as a novel paradigm to adapt LLMs to molecule-caption translation. As shown in Figure 2, ICMA incorporates three stages, including Hybrid Context Retrieval, Post-retrieval Re-ranking, and In-context Molecule Tuning. Specifically, Hybrid Context Retrieval first retrieves $N$ rough examples from the training set $\mathcal{D}$ by calculating the similarity between the current query and the molecule-caption pairs from the training set. After that, Post-retrieval Re-ranking with Random Walk and Sequence Reversal is adopted to obtain $n$ refined examples from the $N$ rough examples. Finally, In-Context Molecule Tuning can be performed to update the parameters of LLMs to learn the molecule-text alignment from refined context examples.
III-A Hybrid Context Retrieval
@@
For the Mol2Cap task, ICMA adopts Molecule Graph Retrieval to better refine the retrieval quality. Previously, MolReGPT [10] utilizes the Morgan Fingerprints (Morgan FTS) and Dice similarity to evaluate the similarity between molecules, which encodes the pre-defined handcraft structures as the embedding of the molecule. However, Morgan FTS are typically based on handcrafted chemical feature extraction, which may have limited capability to capture the comprehensive information of complex molecule structures and properties. On the other side, Graph Neural Networks (GNNs) have been widely used to capture topological structures  [22]. Pre-trained on millions of molecule graphs, GNNs could better understand the molecular structure, providing complete chemical semantics [23]. This makes GNNs a better option for molecule similarity calculation. In ICMA, we adopt a pre-trained GNN encoder to obtain the molecule graph embeddings:
@@
where $\textbf{e}_{m}$ denotes the embedding of the given 2D graph $g_{m}$ of the molecule $m$. Specifically, we adopt Mole-BERT [24] as the GNN encoder.
@@
In the Cap2Mol task, we inherit the BM25 Caption Retrieval [25] from MolReGPT [10] as it focuses on the detail matching of molecule captions, showing competitive performance and is much faster than LLM-based methods like Sentencebert [26]. Specifically, given captions in the test set as query captions $Q_{c}$ and the training set of captions $\mathcal{D}_{c}$, the caption similarity ranking function $\mathcal{R}^{c}$ can be denoted as:
@@
III-B Post-retrieval Re-ranking
Although refined retrieval algorithms could bring better retrieval quality, there are still some problems considering the arrangement of context examples. Thus, we propose Post-retrieval Re-ranking with Random Walk and Sequence Reversal to re-rank the priorities and positions of context examples, thus enhancing the quality of In-Context Molecule Tuning.
III-B1 Random Walk
The molecules ranked top by retrieval algorithms can sometimes share too many overlaps, impairing the informativeness of context examples. In this case, for the context diversity and generalization performance of ICMA, it is necessary to give these less similar examples a chance to be visited. Inspired by the Random Walk mechanism in graph theory, we propose Random Walk as a post-retrieval method to select examples from the top-$N$ retrieved results so that examples with lower rank still have a chance to be selected, which provides more useful information and complements the context diversity. Mathematically, we adopt a dynamic chance for examples with different ranks that gradually decays as the rank moves down. Specifically, for the $j$-th example in the $N$ rough results, where $1\leq j\leq$N, the possibility of skipping $p(j)$ is represented as:
$p(j)=p_{max}*\frac{N-j}{N-1},$ (4)
where $p_{max}$ is the maximum skip probability. It can be seen that $p(j)$ will decay to 0% at the $N$-th example. This guarantees that the sampling stage will not result in an empty selection. Once an example is selected, the subsequent selection process begins from the next example in the sequence, which means that if the $j$-th example is chosen, the next potential selection starts from the $(j+1)$-th example.
Consequently, if the $N$-th example is selected, the Random Walk sampling terminates early, which indicates that we may fail to achieve the intended number of $n$ refined examples. Therefore, we need to make sure that this early-stop condition is scarcely activated so that the integrity of sampling can be guaranteed. Empirically, the skipping probability should not be too large to avoid unstable training. Meanwhile, if $N$ is too large (i.e., $N\gg n$), the retrieval quality will be impaired, while if $N$ is too small (i.e., $N\approx n$), it might hurt the diversity. Therefore, as $n$ is normally less than 5 due to the context length limitation, $N$ is set to 10 as a balanced choice. For simplification, the maximum skip probability $p_{max}$ is then set to $(N-1)\%=9\%$ in this work so that $p(j)$ could be simply written as $(N-j)\%$. In this case, let $n=2$ and $N=10$, there is only a likelihood of $\frac{(N-1)!}{100^{N-1}}=3.6288e^{-13}$ that the early-stop is activated, which is nearly indistinguishable from zero, thus generally maintaining the sampling integrity.
III-B2 Sequence Reversal
Due to the training strategy and the inherent characteristic of natural languages, LLMs have difficulty capturing long-range dependencies or relationships between words that are far apart in the input text, namely the distance dependency. The positions of examples in the context might influence the generation results due to the distance dependency of LLMs [27]. In this case, it is of significance to put the most informative example exactly near the current input query.